You find a zero coupon bond with a par value of 15000 and 15 years to maturity? If yield to maturity on this bond is 5.3, what is the price of the bond? Assume semiannual compounding periods.
Answer #1

Maturity Value valueof ZeroCoupon Bond = = (1 + r)

Since ZCB dont pay coupons but pays bonds par value at maturity.

where r is the yield to maturity i.e. 5.3/2 = 2.65% _ _ _ (Semi-annual)

t is the total compounding periods i.e.  15 years* 2 = 30 periods

valueof ZeroCoupon Bond = 15,000 (1 + 0.0265)30

  - 6844.20

Price of Bond today will be $ 6844.20

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.